A 2-kg mass is held at rest at the top of an incline that make a 400 angle with the ground. The incline is 1.5 m long and the coefficient of kinetic friction µk = 0.2. When the mass is released, slides down the inclined with uniform acceleration.

1
Expert's answer
2022-01-28T08:15:00-0500

The value of the acceleration is


"a=g(\sin\theta+\mu_k\cos\theta)=7.8\text{ m/s}^2."